OpenStudy (anonymous):

How many real number solutions are there to the equation 0= -3x^2+x-4? 0 1 2 3

OpenStudy (gorv):

-3x^2+x-4=0 3x^2-x+4=0 x=\[(1+-\sqrt{1^2-4*4*3})\]

OpenStudy (gorv):

divide by 2*3

OpenStudy (nathan917):

for a quadratic ax2+bx+c=0 use the discriminant Δ=b2−4ac in your question a=-3, b = 1 and c= -4 then 1^2 - 4 x (-3) x (-4) = -47 since Δ<0 no real solutions Rules Δ>0 2 unequal real solutions if Δ is a square number the roots are rational. Δ=0 repeated or equal roots and the senario above

OpenStudy (gorv):

as quantity in root is negative so 0 is the number of real solution

OpenStudy (anonymous):

Thanks